OpenStudy (anonymous):

Amir weighs 42 more pounds than Mike . together, they weigh 370 pounds . how much does Steven weigh ?

OpenStudy (ccswims):

? is there any more information because if this is it there isn't enough to determine

OpenStudy (anonymous):

No theres not anymore @ccswims

OpenStudy (ccswims):

are you saying also combined with stevens weight it is 370?

OpenStudy (anonymous):

Yeah @ccswims

OpenStudy (anonymous):

ok I figured it out

OpenStudy (anonymous):

amir ways 206, and the other person weighs 164

OpenStudy (anonymous):

How did u get that ?

OpenStudy (anonymous):

since 206 is 42 than 164

OpenStudy (anonymous):

okay, so I did 370-42 to get 328

OpenStudy (anonymous):

then I divided 328 to get 164

OpenStudy (anonymous):

which is what Mike ways

OpenStudy (ccswims):

you have to divide it amongest 3 people

OpenStudy (ccswims):

not just 2

OpenStudy (anonymous):

oh, I did not see that, I think you are right @ccswims

OpenStudy (anonymous):

No its 2 people @ccswims

OpenStudy (anonymous):

How is it 3 people ?????!

OpenStudy (anonymous):

because it is asking about Steven

OpenStudy (ccswims):

because they want to know STEVENS weight

OpenStudy (ccswims):

not Mikes, if it was about Mike then yea, it would only be 2 people

OpenStudy (anonymous):

what are the options?

OpenStudy (anonymous):

does it mean all together steven,mike, and amir?

OpenStudy (ccswims):

yea

OpenStudy (anonymous):

I think I can try it and see if I get anything

OpenStudy (anonymous):

No its just asking about steven

OpenStudy (anonymous):

And how much he weighs

OpenStudy (ccswims):

i think i got the answer, but idk because the weight i got for Amir is pretty unrealistic

OpenStudy (ccswims):

i got 165.34 fro Steven

OpenStudy (ccswims):

are there any choices you have for this question?

OpenStudy (ccswims):

like, multiple choices?

OpenStudy (anonymous):

No its just a problem where you show ur work

OpenStudy (ccswims):

ok, well what i did was divide 370 by 3 for the 3 people. i then got 123.3. 123.33 - 42 = 81.33. 123.33 + 83.33 = 204.66 370 - 204.66 = 165.34

OpenStudy (ccswims):

so basically i got Amir = 84.33 Mike = 123.33 Steven = 165.34

OpenStudy (ccswims):

but idk because Amir's weight seems unrealistic

OpenStudy (anonymous):

I don't think that is right because when I added them I didn't get 370

OpenStudy (anonymous):

Me either

OpenStudy (ccswims):

WAIT

OpenStudy (ccswims):

i meant 81.33 not 84.3

OpenStudy (ccswims):

but now thats even more unrealistic

OpenStudy (anonymous):

I was subtracting 370-42 then got 328 should I divide 328÷3 ??

OpenStudy (anonymous):

@ccswims That seems about right, but I don't think the problem is very realistic like you wrote/said

OpenStudy (anonymous):

I don't think there is a problem with your answer but rather a problem with the problem XD

OpenStudy (anonymous):

Yeah this problem was very stupid !!!